Again, please use the same data as in the previous question. It will be repeated for your convenience each time.

The following is production data:

Year Waffles
(Quantity)
Waffles
(Price)
Orange Juice
(Quantity)
Orange Juice
(Price)
Toaster Ovens
(Quantity)
Toaster Ovens
(Price)
2010 1000 $1.20 400 $5 280 $22
2011 1200 $1.30 410 $6 300 $25
2012 1500 $1.40 420 $7 300 $28

The above, like I said, is production data. Consumption data is as follows: the average citizen consumes 80 waffles, 20 units of orange juice and 10 toaster ovens. 2010 is the base year for all calculations.

Carefully following all directions for entering your answer, calculate the inflation rate, using the GDP deflator, for 2010-2011.

Express your answer "as a percent, but without the percentage sign." (See the math review if you need clarification.) In other words, if your answer is 99.99 percent (0.9999 in decimal format), enter only 99.99 in the blank. Make sure to round all intermediate steps to four decimal places and your final answer to two. In this case, the final step will be conversion into percent.
Answer #1

Answer : for 2011 :

Nominal GDP = Current year price level * Current year production level

In 2011 the nominal GDP = ($1.30 * 1200) + ($6 * 410) + ($25 * 300) = 1,560 + 2,460 + 7,500 = $11,520

Real GDP = Base year price level * Current year production level  

Here Base year is 2010.

In 2011 the real GDP = ($1.20 * 1200) + ($5 * 410) + ($22 * 300) = 1,440 + 2,050 + 6,600 = $10,090

GDP deflator = (Nominal GDP / Real GDP) * 100

In 2011 the GDP deflator = (11,520 / 10,090) * 100 = 1.1417 * 100

=> In 2011 the GDP deflator = 114.17

Therefore, here in 2011 the GDP deflator is 114.17 .

For 2010 :

Nominal GDP of 2010 = ($1.20 * 1000) + ($5 * 400) + ($22 * 280) = 1,200 + 2,000 + 6,160 = $9,360

Real GDP of 2010 = ($1.20 * 1000) + ($5 * 400) + ($22 * 280) = 1,200 + 2,000 + 6,160 = $9,360 [As 2010 is the base year].

GDP deflator = (Nominal GDP / Real GDP) * 100

In 2010 the GDP deflator = (9,360 / 9,360) * 100

=> In 2010 the GDP deflator = 100

Inflation rate = [(GDP deflator of 2011 - GDP deflator of 2010) / GDP deflator of 2010] * 100

=> Inflation rate = [(114.17 - 100) / 100] * 100

=> Inflation rate = 14.17

Therefore, here the inflation rate between 2010 and 2011 is 14.17 .